OpenStudy (anonymous):

solve 2cos3(x-pi/4)=0

OpenStudy (anonymous):

same as \(\cos(3(x-\frac{\pi}{4}))=0\)

OpenStudy (anonymous):

use the "equation" option to make it luk understandable

OpenStudy (anonymous):

for xE[-2pi,2pi]

OpenStudy (anonymous):

\(\cos(\frac{\pi}{2})=0\) so for one solution you can solve \(3(x-\frac{\pi}{4})=\frac{\pi}{2}\) for \(x\)

OpenStudy (anonymous):

\[2\cos3(90-(135-x) )\]

OpenStudy (anonymous):

kk \[2\cos3(x-\frac{\pi}{4})=0 for -2\pi \le x \le 2\pi\]

OpenStudy (anonymous):

4 cos^3 (x) - 3cos x= cos3x

OpenStudy (anonymous):

@satellite73 um how to get rest of solutions within domain

OpenStudy (anonymous):

im missing a few solns

OpenStudy (anonymous):

well, you could keep solving i guess

OpenStudy (anonymous):

probably a quicker way to do it, but you have \(\cos(\frac{3\pi}{2})=0\) so you could solve again

OpenStudy (anonymous):

let me think if there is a snap way to find the others probably is

OpenStudy (anonymous):

lets do it like adults we know that \(\cos(\frac{(2n+1)\pi}{2})=0\) so we want to solve \[3(x-\frac{1}{4})=\frac{(2n+1)\pi}{2}\] for \(x\)

OpenStudy (anonymous):

sorry, last line should read \[3(x-\frac{\pi}{4})=\frac{(2n+1)\pi}{2}\]

OpenStudy (anonymous):

is that just the general soln

OpenStudy (anonymous):

cant you just solve this like you would any cos function ?

OpenStudy (anonymous):

\[3x-\frac{3\pi}{4}=\frac{(2n+1)\pi}{2}\] \[3x=\frac{(2n+1)\pi}{2}+\frac{3\pi}{4}\] \[3x=\frac{4n\pi+2\pi+3\pi}{4}\] \[3x=\frac{(4n+1)\pi}{4}\] \[x=\frac{(4n+1)\pi}{12}\]

OpenStudy (anonymous):

um so what do i do with that